How many times do you need to toss a fair coin to get 100 heads with probability 90%?


1
Expert's answer
2021-05-07T14:25:25-0400

Let we toss a coin n times-


Probability of getting 1 heads "= ^nC_1(\dfrac{1}{2})^1)\dfrac{1}{2})^{n-1}"

"= n\dfrac{1}{2^n}"


So Probability of getting 100 head "=\dfrac{100n}{2^n}"


As per Question-


"\dfrac{100n}{2^n}=0.9\\[9pt]

100n=9.2^n"


At n=9, The condition is satisfied.


So, The number of times we have to toss is 9.